# LLM Critics Help Catch Bugs in Mathematics: Towards a Better Mathematical Verifier with Natural Language Feedback

### v3

@@ 5.4 Ablation Study
Removing stage 1 essentially reverts our method to a vanilla ORM, as shown in the table. Without training on natural language feedback, the model is unaware of the reasons behind what makes an answer correct or incorrect. Hence, the performance of the binary classification in the second stage noticeably declines compared to MATH-Minos.
When eliminating stage 2, binary classification, an intuitive approach is to directly utilize the natural language feedback generated by the text reward model for the generator’s verification. Given that the model outputs a binary discrete value (’True’ or ’False’), we cannot employ a best-of-N verification but instead use it as a filter. Specifically, we apply self-consistency in filtering out cases where the model output is ’True’. Unfortunately, we observe that relying solely on natural language feedback from text reward model leads to a significant decline in performance. The probable reasons may include: 1) Upon closer inspection, we notice inconsistencies in the model’s feedback. This is characterized by samples where a step is recognized as incorrect yet the overall outcome is deemed correct, and vice versa. Such inconsistencies are even found in the strongest models such as GPT-4, despite their ability to provide accurate explanations. 2) The performance of evaluation might be constrained by the model’s scale. Influenced by computational resources, we do not further explore larger models. Evaluation generation tasks could be more challenging for models of smaller scale. 3) The binary discrete output of the model is relatively coarse-grained. For instance, two examples judged as correct cannot be compared with each other.
@@
5.5 Performance on False Positive Samples
@@
To further investigate the efficacy of Math-Minos, we analyze the performance on false positive samples within the training dataset. False positive samples refer to those instances that have a correct final outcome but contain errors in the intermediate steps. Ideally, a robust verifier should assign lower rewards to these samples. We extract such examples from the training set of the verifier, amounting to a total of 600 samples, which includes data from both GSM8K and MATH datasets. We test the performance of both ORM and Math-Minos, with the experimental results presented in Table 5.
According to our experimental findings, it turns out that vanilla ORM can correctly discriminate a majority of the false-positive samples from the training set but with an accuracy significantly lower than MATH-Minos. It achieves a recall of 74% with an average reward of 0.234. While MATH-Minos reach a recall rate of 92% with an average reward of 0.105. This performance is significantly better than that of ORM not trained on the natural language feedback. Delving into the data, we discover that in the context of false positives, a substantial portion of the natural language feedback generated by GPT-4 are contradicted to the “True” labels we assigned. We believe that these data endows MATH-Minos with a stronger capability to discern false-positives, thereby enhancing the model’s performance.
5.6 Influence of the Data Amount
@@
Figure 6 illustrates how different amounts of natural language feedback affect the model during the first stage. We use SFT on the model in the first stage using different scales of natural language feedback. In the second stage, we adopt the setup of ORM setting and use the verifier to select the best-of-N of GSM8K test set. We observe a positive correlation between the model’s performance and the quantity of natural language feedback provided in stage one, which implicitly evidences the benefit of natural language feedback for the model.
6 Conclusion
We analyze the current training paradigm of verifiers, demonstrating that score feedback from binary classification labels is suboptimal for teaching LLMs to accurately evaluate mathematical solutions. By introducing rationale labels that provide detailed explanations of error types, our training paradigm significantly enhances the verifier’s evaluation ability. The experimental results show that models trained on a small dataset with natural language feedback (30k instances) significantly outperform the baselines that rely solely on classification labels. This highlights the critical role of rich and informative labels in training data in crafting more nuanced and effective training strategies for the development of large language models (LLMs) that are capable of complex reasoning tasks. Finally, we hope the findings of this work can pave the way for the potential integration of natural language feedback with classification verifiers.
7 Limitations
Following the scaling laws, the evaluation ability of a model, especially in terms of generating natural language evaluations, may vary across different sizes. However, due to computational resource limitations, experiments were conducted solely on a model with 7 billion parameters, thereby unable to explore the impact of model’s scaling on the evaluation ability. We leave it into our future work.

### v4

@@ 5.4 Ablation Study
Removing stage 1 essentially reverts our method to a vanilla ORM, where the model is unaware of the reasons behind what makes an answer correct or incorrect. Hence, the performance noticeably declines compared to MATH-Minos.
When eliminating stage 2, binary classification, an intuitive approach is to directly utilize the natural language feedback generated by the text reward model for the generator’s verification. Given that the model outputs a binary discrete value (’True’ or ’False’), we cannot employ a best-of-N verification but instead use it as a filter. Specifically, we apply self-consistency in filtering out cases where the model output is ’True’. Unfortunately, we observe that relying solely on natural language feedback from the textual reward model leads to a significant decline in performance. The probable reasons may include: 1) Upon closer inspection, we notice inconsistencies in the model’s feedback. This is characterized by samples where a step is recognized as incorrect yet the overall outcome is deemed correct, and vice versa. Such inconsistencies are even found in the strongest models such as GPT-4, despite their ability to provide accurate explanations. 2) The performance of evaluation might be constrained by the model’s scale. Influenced by computational resources, we do not further explore larger models. Evaluation generation tasks could be more challenging for models of smaller scale. 3) The binary discrete output of the model is relatively coarse-grained. For instance, two examples judged as correct cannot be compared with each other.
@@
6 Conclusion
We investigate the effectiveness of the current training paradigm of verifiers, demonstrating that score feedback from binary classification labels is suboptimal for teaching LLMs to accurately evaluate mathematical solutions. By introducing natural language feedback with a two-stage training paradigm, we significantly enhance the verifier’s evaluation ability. The experimental results show that models trained on a small dataset with natural language feedback (30k instances) significantly outperform the baselines that rely solely on classification labels. This highlights the critical role of rich and informative labels in training data in crafting more nuanced and effective training strategies for the development of LLMs that are capable of complex reasoning tasks. Finally, the findings of this work pave the way for the potential integration of natural language feedback with classification verifiers.
7 Limitations
Following the scaling laws, the evaluation ability of a model, especially in terms of generating natural language evaluations, may vary across different sizes. However, due to computational resource limitations, experiments were conducted solely on a model with 7 billion parameters, thereby being unable to explore the impact of the model’s scaling on the evaluation ability. We leave it to our future work. We do not verify the Math-Minos (PRM) in the reinforcement learning setting because of the lack of open-source implementation of step-by-step PPO. We also leave it to our future work.
Appendix A Performance on False Positive Samples
@@
To further investigate the efficacy of Math-Minos, we analyze the performance on false positive samples within the training dataset. False positive samples refer to those instances that have a correct final outcome but contain errors in the intermediate steps. Ideally, a robust verifier should assign lower rewards to these samples. We extract such examples using the step-level labels from the training set of the verifier, amounting to a total of 600 samples, which includes data from both GSM8K and MATH datasets. We test the performance of both ORM and Math-Minos, with the experimental results presented in Table 7.
According to our experimental findings, it turns out that vanilla ORM can correctly discriminate a majority of the false-positive samples from the training set but with an accuracy significantly lower than MATH-Minos. It achieves a recall of 74% with an average reward of 0.234. While MATH-Minos reach a recall rate of 92% with an average reward of 0.105. This performance is significantly better than that of ORM not trained on the natural language feedback.
Delving into the data, we discover that in the context of false positives, a substantial portion of the natural language feedback generated by GPT-4 are contradicted to the “True” labels we assigned. We believe that these data endows MATH-Minos with a stronger capability to discern false-positives, thereby enhancing the model’s performance.
Appendix B Influence of the Data Amount
@@
Figure 7 illustrates how different amounts of natural language feedback affect the performance of MATH-Minos during the first stage. We use SFT on the model in the first stage using different scales of natural language feedback. In the second stage, we adopt the setup of ORM setting and use the verifier to select the best-of-N of GSM8K test set. We observe a positive correlation between the model’s performance and the quantity of natural language feedback provided in stage one, which implicitly evidences the benefit of natural language feedback for the model.
Appendix C Comparison with COT distillation
Model GSM8K MATH500
ORM 86.2 35.9
Math-Minos(ORM) 87.3 37.4
ORM w/ GPT-4-COT-Distillation 85.8 36.2
Table 8: Experimental result of different ways incorporating GPT-4 chain-of-thought data. "ORM w/ GPT4-COT-Distillation" denotes we use two-stage training of ORM with direct COT distillation generated by GPT-4.
To further validate the efficacy of our method, we compared it against a more generalized baseline approach under conditions where the same number of GPT-4 tokens is utilized. Specifically, we employed enhanced queries from Meta-MATH and consequently distilled the problem-solving processes generated by GPT-4 into the ORM training. We aimed to determine whether training the model with COT data distilled using the same number of GPT-4 tokens could yield comparable results with MATH-Minos.
The experimental results are as follows: Our findings indicate that using an equivalent number of GPT-4 tokens on COT-distillation resulted in negligible improvements to the verifier. This lack of enhancement can be attributed to the fact that its base model, Mistral-7B:MetaMATH, has already been trained on a substantial amount of high-quality COT data. Consequently, this step contributed almost no additional informational gain to the model. These results further underscore the effectiveness and necessity of our Label-Aware Natural Language Feedback Curation.
Appendix D Detailed Experimental Setup
We use the MetaMATH-Mistral as the generator for the questions in the test set. In order to ensure the model has the ability to solve mathematical problems before learning to evaluate, we also use the MetaMATH-Mistral as the base model for MATH-Minos and all other reward models. For the training of natural language feedback, we use 30k training data generated as Section 3.2 with a learning rate of 5e-6 with a total batch size of 256. For the training of score feedback, we use 440k training data (i.e., 30k data with the binary classification labels from the training data in the training of natural language feedback and 410k data sampled from MATH-Shepherd. For the training of baseline, we use the total 440k training data from MATH-Shepherd. For the training of ORM, we adopt the learning rate of 3e-6 with a batch size of 512. For PRM, the learning rate is 2e-6 with a batch size of 512. In reinforcement learning, we follow the setting of Wang et al. (2024) with the learning rate of 1e-7.
Appendix E Case Study
The Figure 8 illustrates the different content generated by GPT-4-turbo between Label-Aware Natural Language Feedback Curation and direct prompts. Our proposed method, guided by labels, helps the model mitigate the impact of hallucinations to some extent, resulting in high-quality natural language evaluation.
Figure 8: Case study of the Label-Aware Natural Language Feedback Curation.
Figure 9: The detailed prompt template used for prompting GPT-4-turbo to generate Label-Aware Natural Language Feedback.
Figure 10: The detailed prompt template used for prompting GPT-4-turbo to generate the error analysis reason.